The short-run total cost function of a firm that employs labour (L) and fixed capital is given by c = vKo + wq^1/B × Ko^ -a/B Where w is the cost of the labour and v is the cost of capital (I) Derived the marginal cost of the firm ii. Assuming that the firm is a price taking one that sells its output at p per unit, derive the short-run supply function of the firm iii. If there are 200 firms in the industry with similar cost conditions, compute the total market supply.
1
Expert's answer
2020-04-03T10:04:36-0400

(i)"c=K0\times (v+w\times q^\frac{-a}{b^2})"

derivative function

"c=v+w\times q^\frac{-a}{b^2}"

the marginal cost of the firm

(ii) supply volume will be determined by the point of intersection of the price line with the marginal cost curve:

"p=v+w\times q^\frac{-a}{b^2}"

(iii) "QS=c+d\times p"

"QS=200\times (v+w\times q^\frac{-a}{b^2})+d\times p"